OpenStudy (anonymous):

How do I show the proof for: For all integers a, if a^2 = 0(mod 6), then a = 0(mod 6)? I think the answer is true, but I don't know how to prove it generally.

OpenStudy (anonymous):

a=0(mod 6) multiplying both sides by a give us : a^2=0(mod 6)

ganeshie8 (ganeshie8):

thats the converse direction, but we wanto prove this : given a^2=0 mod 6, prove a = 0 mod 6

OpenStudy (loser66):

let me try :) @ganeshie8 watch me, please :) a^2 =0 (mod 6) a^2-0 =0 (mod6) (a+0)(a-0) =0 (mod 6) a+0 =0 (mod6) a =0 (mod 6)

OpenStudy (anonymous):

:)

ganeshie8 (ganeshie8):

looks legit :) u have used quadratic congruences !

ganeshie8 (ganeshie8):

another long/boring way is : a^2 = 0 mod 6 => a^2 = 0 mod 2 a^2 = 0 mod 3 2 | a^2 => 2 | a 3 | a^2 => 3 | a since gcd(2,3) = 1, 2*3 | a a = 0 mod 6

OpenStudy (loser66):

hihihi....

Parth (parthkohli):

Like you did 2 | a^2 => 2 | a, why not 6 | a^2 => 6 | a?

ganeshie8 (ganeshie8):

6 is not prime

Parth (parthkohli):

OK, I got it!

ganeshie8 (ganeshie8):

im looking for a contradiction to show n | a^2 does not necessarily mean n | a but not getting any quick examples...

ganeshie8 (ganeshie8):

n is composite..

ganeshie8 (ganeshie8):

a = b (mod n) division is not allowed here unless 'a' and 'n' are coprime... so it shouldnt work in general....

OpenStudy (amistre64):

0 1 2 3 4 5 0 0 0 0 0 0 0 1 0 1 2 3 4 5 2 0 2 4 0 2 4 3 0 3 0 3 0 3 4 0 4 2 0 4 2 5 0 5 4 3 2 1 0^2 = 0 1^2 = 1 2^2 = 4 3^2 = 3 4^2 = 4 5^2 = 1 proof by table

OpenStudy (anonymous):

:)

OpenStudy (loser66):

learn every day!!

OpenStudy (anonymous):

@ganeshie8 you said division is not allowed, so how do you consider the Quadratic congruence as you call it accepted ?

ganeshie8 (ganeshie8):

direct division is not allowed, division is defined as below :- a = b mod n a/c = b/c mod (n/gcd(c, n) )

ganeshie8 (ganeshie8):

when u divide, u need to divide the mod also by gcd(c, n) oly dividing the remainders will not work

OpenStudy (anonymous):

(a+0)(a-0) =0 (mod 6) to a+0 =0 (mod6) isn't a direct division ?

ganeshie8 (ganeshie8):

thats just zero product property which is legal

ganeshie8 (ganeshie8):

it is legal in congruences also, not just algebra..

ganeshie8 (ganeshie8):

ab = 0 => either one of them or both are 0

OpenStudy (anonymous):

Ok

ganeshie8 (ganeshie8):

here is one contradicting example for the statement : a^2 = 0 mod n does not necessarily mean a = 0 mod n 4^2 = 0 mod 16 , but 4 = 0 mod 16 is false.

OpenStudy (anonymous):

isn't mod 6

ganeshie8 (ganeshie8):

we're talking in general

ganeshie8 (ganeshie8):

not specific to a 6 or something... 6 is not a special number... not worth remembering its properties by heart :)

OpenStudy (anonymous):

good, How do you prove that 1^n+2^n+3^n+4^n=0 mod 5 ?

ganeshie8 (ganeshie8):

easy : 4 = -1 3 = -2 everything cancel out

ganeshie8 (ganeshie8):

u doing number theory or groups ?

OpenStudy (anonymous):

or 1 and 2 congruent to -3 and -4 right ?

ganeshie8 (ganeshie8):

yes, 1 = -4 2 = -3

OpenStudy (anonymous):

so it means always the number of sum parts must be even ?

ganeshie8 (ganeshie8):

im not sure we can generalize like that... wat do we get by generalizing that anyway ?

ganeshie8 (ganeshie8):

try this : find the remainder when \(\large 86^{1234921897129}\) is divided by 87

ganeshie8 (ganeshie8):

yup ! too easy for u Loser :)

ganeshie8 (ganeshie8):

nobody can help u cuz u are doing all tough problems nobody ever messed wid lol